OpenStudy (anonymous):

An experiment consists of flipping a fair coin 3 times. What are the odds in favor of obtaining at least one head?

OpenStudy (anonymous):

1/2*1/2*1/2 =1/6, 6/6-1/6=5/6 I think that's right